OpenStudy (anonymous):

(x + 3)2/3 + 2(x + 3)1/3 - 15 = 0

OpenStudy (lalaly):

\[\frac{2}{3}x+2 + \frac{2}{3}x+2-15 =0\]\[\frac{4}{3}x=11\]\[4x=33\]\[x=\frac{33}{4}\]

jhonyy9 (jhonyy9):

4(x+3)/3 =15 4(x+3)=45 4x=33 x=33/4